Spud, Inc. a manufacturer of gourmet potato chips, employs activity-based costing. The budgeted data for each of the activity cost pools is provided below for the year 2017.

Activity Cost Pools

Estimated
Overhead

Estimated Use of
Cost Drivers per Activity

Ordering and receiving $101,568 13,800 orders
Food processing 483,125 62,500 machine hours
Packaging 1,783,500 435,000 labor hours


For 2017, the company had 11,000 orders and used 50,000 machine hours, and labor hours totaled 502,000.

(a)

Calculate the overhead rates for each activity. (Round answers to 2 decimal places, e.g. 12.25.)

Overhead Rates

Ordering and receiving

$

per order
Food processing

$

per machine hour
Packaging

$

per labor hour

b) What is the total overhead applied?
Answer #1

Overhead rate=Respective cost/Respective activity

Overhead Rates
Ordering and receiving (101,568/13800)=$7.36 per order
Food processing (483125/62500)=$7.73 per machine hour
Packaging (1,783,500/435,000)=$4.1 per labor hour

Total overhead applied=Overhead rate*Respective activity

=(7.36*11000)+(7.73*50,000)+(4.1*502000)

=$2525660